It is required to inclose a rectangular field by a fence and then to divide it into two lots by a fence parallel to one of the sides. If the area of the field is given, find the ratio of the sides so that the total length of fence shall be a minimum.


1
Expert's answer
2022-05-30T16:19:52-0400

Let "x=" the length of the field, let "y=" the width of the field.

Given "A=xy=const."

Then "y=\dfrac{A}{x}."

The total length "P" of fence will be


"P=2x+2y+y"

Substitute


"P=P(x)=2x+3(\dfrac{A}{x})"

Differentiate with respect to "x"


"P'=2-\dfrac{3A}{x^2}, x>0"

Find the critical number(s)


"P'=0=>2-\dfrac{3A}{x^2}=0"

"x=\pm \sqrt{\dfrac{3A}{2}}"

Since "x>0" we take "x=\sqrt{\dfrac{3A}{2}}."

If "0<x<\sqrt{\dfrac{3A}{2}}, P'<0, P(x)" decreases.

If "x>\sqrt{\dfrac{3A}{2}}, P'>0,P(x)" increases.

The function "P(x)" has a local minimum at "x=\sqrt{\dfrac{3A}{2}}."

Since the function "P(x)" has the only extremum for "x>0," then the function "P(x)" has the absolute minimum for "x>0" at "x=\sqrt{\dfrac{3A}{2}}."


"y=\dfrac{A}{\sqrt{\dfrac{3A}{2}}}=\sqrt{\dfrac{2A}{3}}"

"length/width=\dfrac{x}{y}=\dfrac{\sqrt{\dfrac{3A}{2}}}{\sqrt{\dfrac{2A}{3}}}=\dfrac{3}{2}"

if the fence is parallel to the width of the field then the ratio "length\/width" is "3\/2."